The width of a rectangle is 6 centimeters less than the length.The perimeter is
64 centimeters.Find the width.

Answer by checkley77(12844)

W=L-6
P=2L+2W
64=2L+2(L-6)
64=2L+2L-12
64=4L-12
4L=64+12
4L=76
L=76/4
L=19 ANS. FOR THE LENGTH.
W=19-6=13 ANS. FOR THE WIDTH.
PROOF:
64=2*19+2*13
64=38+26
64=64